A particle perform SHM of period 12 second and amplitude 8 cm .if initially the particle is at positive extremity, how much time will it take to cover a distance of 6 cm from the extreme position

Grade:12

1 Answers

Amritpal Singh Saini
5 years ago
At extreme position the displacement is given by X=acos(at) , X,a, are given in the question as 6 and 8 respectively, also w=2pi/T (put value of T)THUS , w=pi/6 ..... On substitution of all the values to the main equation and solving we would get answer as 0.86 sec
The equation is X=acos(wt) where w=2pi/ T , thus w=pi/6 , X/a =( cos pi/6)t Thus t= 0.86 sec(It was a typing error for the equation)